OpenStudy (anonymous):

f(x)=-6x-1 and g(x)=4x^2 find the solution for (fg)(x)

OpenStudy (zzr0ck3r):

\((fg)(x)=f(x)*g(x) = (-6x-1)(4x^2)\)

OpenStudy (anonymous):

then you distribute?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

or leave it alone... either way.

OpenStudy (anonymous):

then if I distribute it becomes -24x^2-1?

OpenStudy (zzr0ck3r):

\(4x^2(-6x-1)=(4x^2)(-6x)-(4x^2)*1=-24x^3-4x^2\)

OpenStudy (anonymous):

are you positive because that doesn't seem like what I got the other day

OpenStudy (zzr0ck3r):

Dont know what to tell you.

OpenStudy (zzr0ck3r):

http://www.wolframalpha.com/input/?i=%28%E2%88%926x%E2%88%921%29%284x2%29 scroll down to where it says "expanded form".

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how about 5/8+sqrt 7 I couldn't enter this in